OpenStudy (anonymous):

solve each equation for y. If the equation is a direct variation, identify the constant of variation. 7. y+4=6x+4 8. 12+y=3x 9. y/x =12

OpenStudy (anonymous):

y + 4 = 6x + 4 implies y = 6x + 4 - 4 This gives y = 6x which is a direct variation with 6 as the constant of variation

OpenStudy (anonymous):

In order words, direct variations are in the form y = kx where k = constant of variation.